# Covariance spectrum of MAXI J1820+070: On the nature of the Comptonizing flow

### v1

@@ 3.1.3 The root mean square and covariance spectra
The time-averaged X-ray spectra alone do not capture how the individual spectral components vary relative to each other on timescales of seconds to minutes. Instead, the PDS and coherence analyses reveal significant differences in variability across various energies and frequencies (Paice2021; Tetarenko2019; Tetarenko2021). To further investigate the properties of the variability and their correlations, energy-dependent spectral-timing analyses are essential.
@@
Wilkinson2009 developed a technique known as the covariance spectrum. This covariance technique can be considered a matched filter applied to the data. By utilizing the PDS of both the reference band and the band of interest, we can derive the cross-spectrum and subsequently obtain a Fourier frequency-resolved covariance spectrum. This approach allows us to identify the variability in the energy channel being examined that is correlated with the reference band (Uttley2014). For further details, please refer to the Appendix.
@@
We can obtain the RMS spectrum by integrating the PDS over a specified frequency range. The RMS is background-corrected using the formula $RMS=\sqrt{P}(S+B)/S$, where S and B represent the source and background count rates, respectively, and P is the power calculated over the selected frequency (Bu2015). The RMS spectrum illustrates the overall intrinsic variability, while the covariance spectrum emphasizes the variability that correlates with the reference band. If there is a single variability pattern across the entire energy band and coherence is perfect (unity), the shape of the covariance spectrum will match that of the RMS spectrum (Wilkinson2009).
@@
Figure 4: Fractional covariance spectra. The blue and red points represent the spectrum of long and short-timescale noise, respectively. The reference band is selected as 2–10 keV.
3.1.4 Fractional covariance spectra
@@
By dividing the covariance spectra by the time-averaged spectra, we derive the fractional covariance spectra. Similar to the fractional RMS spectra, the fractional covariance spectra allow us to estimate the contribution of each component to the overall variability (Cassatella2012).
@@
Figure 4 displays the fractional covariance spectra obtained on both long and short timescales, using the 2-10 keV reference band. At low energies, the fractional RMS spectra and the fractional covariance spectra closely align, appearing almost flat. And the coherence approaches unity. In contrast, at high energies, the fractional covariance spectra exhibit a power-law slope. Notably, the fractional covariance spectrum over the long timescale is significantly harder than that over the short timescale.
3.1.5 Covariance ratio
@@
We calculate covariance ratios by dividing the long-timescale covariance spectrum by the short-timescale covariance spectrum. By utilizing more than 70 observations from Insight-HXMT during the X-ray hard state, we investigated the evolution of the covariance ratio throughout the outburst. As mentioned earlier, due to the enormity of the dataset, here we only display the covariance ratios for the five representative epochs of observations in Figure 5. The covariance ratios remain relatively constant with energy below approximately 30 keV. However, above 30 keV, the covariance ratios increase with energy, indicating a hard excess of variability on long timescales. In other words, the trend of coherence toward higher energies on long timescales differs from that on short timescales. This suggests that the variability in the hard band exhibits distinct behaviors on both long and short timescales. An increase in the covariance ratio may indicate additional Comptonization variability on long timescales.
As illustrated in Figure 5, the covariance ratio on MJD 58203 displays a significantly steeper slope than in adjacent observations. This indicates that the profile of the covariance ratio evolves over time. In the first epoch, the covariance ratio increases gradually with increasing energy. In subsequent epochs, the trend of the covariance ratio increasing with energy becomes more pronounced. Notably, a hard excess appears abruptly above approximately 30 keV on long timescales. As previously discussed, the hard X-ray component, which correlates with the soft X-ray component, spans a broad energy range across all timescales but shows greater variation on longer timescales.
By analyzing the significant changes in the covariance ratio with energy, we can see in a model-independent way that the variability of the hard band (>> 30 keV) emission, which is linearly correlated with the variability of the soft band (2-10 keV) emission, exhibits distinct behaviors on both long and short timescales.
@@
3.2 Modeling of the covariance spectra
@@
In this study, we examined the covariance spectra, which are related to spectral variability. We have observed a significant hard excess of the variability over long timescales. To identify the origin of the hard excess, we will fit the long- and short-timescale covariance spectra.
@@
For the covariance spectra, we employed a simple model Tbabs*(diskbb+Nthcomp), since the reflection features are not required in the fits. For each observation, we simultaneously fit the short- and long-timescale covariance spectra while constraining the absorbing column density of Tbabs to $N_{\mathrm{H}}=1.5\times 10^{21}\rm cm^{-2}$, which corresponds to the value used in the fits of the time-averaged energy spectra (you2021). The remaining parameters were allowed to vary freely between the covariance spectra. 72 of 76 covariance spectra can be fitted well with the reduced $\chi^{2}<1.0$.
The best-fitting covariance spectra and the time-averaged spectrum are presented in Figure 6. Part of the fitted parameters for energy and covariance spectra are listed in Table 2. The covariance spectra on different timescales are consistent at low energies but show different electron temperatures, which implies that the hard excess is not associated with the soft component, but instead arises entirely from the hard X-ray emission.
Assuming that the variability of the Compton component mainly results from changes in its normalization, the covariance spectrum in its dominant energy band would closely mirror the shape of the time-averaged energy spectrum. The best-fitting parameters for the short-timescale covariance spectrum are generally consistent with those from the long-timescale covariance spectrum, except for the electron temperature. The electron temperature measured on the long timescale is significantly higher than that measured over the short timescale.
@@
Additionally, we analyze the evolution of electron temperature over time. As shown in Figure 7, early in the hard state, electron temperature on the long timescale is indeed greater than that on the short timescale. Interestingly, electron temperature on the long timescale remains relatively constant throughout the hard state. In contrast, the electron temperature on the short timescale exhibits an early decrease, followed by a distinct increase, which is anticorrelated with the X-ray luminosity during the hard state.
Figure 7: The evolution of the electron temperature in the model Nthcomp is estimated from fits of covariance spectra over long timescales (red dots) and short timescales (blue dots).
4 Discussion
In this study, we study energy-dependent coherence up to 150 keV and discover that the hard X-ray variability above approximately 30 keV is incoherent with the soft X-ray variability (2-10 keV) (see Fig. 3). This finding suggests that at least two distinct Comptonization components contribute to the overall X-ray variability. One component varies linearly with the soft X-ray variability and primarily emits below 30 keV. The other component varies incoherently with the soft X-ray variability and primarily emits above 30 keV. Hereafter, we refer to these components as the coherent component and the incoherent component, respectively.
@@ 4 Discussion
Furthermore, we fit the covariance spectra for both timescales, using the model Tbabs*(diskbb+Nthcomp), to quantitatively characterize the covariance spectra (see Fig. 5). The parameters obtained from the short-timescale covariance spectrum are generally consistent with those from the long-timescale covariance spectrum, except for the electron temperature. The evolution of electron temperature differs notably between the two timescales (see Fig. 7). The electron temperature associated with long-timescale variability is higher than that associated with short-timescale variability, indicating a deficit in hard X-ray variability on the short timescale. Specifically, the electron temperature for the long-timescale covariance spectrum remains relatively constant throughout the hard state, while the short-timescale electron temperature initially decreases before exhibiting a distinct increase. Thus, the evolution of the electron temperature for short timescale variability directly influences the evolution of the covariance ratio. This suggests that the increase in covariance ratios above approximately 30 keV is due to a deficit of hard X-ray variability on short timescales.
4.1 On the incoherent variability: its origin
@@
Since Comptonization emission is influenced by both seed photons and electrons, the incoherent variability is likely due to the complexity of these two components. In this section, we will discuss the potential regions where Comptonization occurs and the origins of the seed photons in order to better understand the nature of this variability. Furthermore, we will explore the association between the Comptonizing source and the flows near the BH.
4.1.1 Multiple Comptonization emission
@@
One possibility accounting for the observed incoherent variability is that multiple Comptonization emissions, as proposed in past studies, are involved. For instance, Dzielak2021 studied the BBN of the PDS for MAXI J1820+070. They found that the BBN could be modeled with four Lorentzian components, each exhibiting significant variability. At least two Comptonization components, characterized by different electron temperatures and optical depths, are required to accurately fit both the variability and time-averaged spectra. Zdziarski_2021a studied the broadband spectra and found that a multi-Comptonization model can fit the spectra well. In their scenario, the system consists of two distinct Comptonization components: a harder component originating from a larger-scale height accretion flow located downstream of the truncation radius, and a softer component forming a corona over the inner disk. These separate regions produce distinct Compton reflection signatures. Wang2024 examined the evolution of the frequency-dependent RMS-flux relation in MAXI J1820+070 using Insight-HXMT data. The timing features and their evolution support the existence of two distinct Comptonization regions: one that remains radially extended, allowing fluctuations to propagate from the disk to the corona, and another that is vertically extended and continues to expand during the later stages of the initial hard state.
@@
It should be noted that we cannot completely rule out a single source of Comptonization. The exponential cut-off function is highly nonlinear, meaning that variations in the cut-off energy can reduce coherence. If these variations exist, the resulting non-linear changes could cause the covariance spectrum to differ from the time-averaged spectrum, as it would be influenced by the cut-off variations. It seems plausible that the incoherent variability could be partially attributed to this type of non-linear spectral variability. However, to draw more definitive conclusions, we would need further quantitative modeling that might suggest the presence of distinctly varying components beyond mere non-linearity.
4.1.2 Multiple fields of seed photon
@@
Another possibility responsible for the incoherent variability is associated with the complexity of the seed photon field. To examine this, we first consider disk photons as the seed photons involved in the inverse Compton scattering process. Uttley2025 constructs a model in which the interplay between the disk seed photons and the corona is taken into account to simulate the Comptonization variability. Based on their model, we conducted a simulation considering 10 groups of disk photons with different seed energies. All of these photons were Compton up-scattered by thermal electrons in the corona via inverse-Compton processes. Since the broadband noise arises due to propagating accretion rate fluctuations, the input seed photon variations should be almost the same. In this case, we found that we could not reproduce the reduction of the coherence in the hard X-ray band (see Fig. 3), as all the resulting power-law emission variations remained coherent. We could reproduce only the broad reduction in coherence when using independent input signals for each power-law component.
@@
In this framework, we conclude that the seed photons for inverse-Compton scattering must have two origins: the disk blackbody and synchrotron emission (yang2015), and it is the Comptonization of synchrotron emission as seed photons that is responsible for the incoherent variability.
4.2 On the coherent variability: electron temperature on different timescales
@@
Variability on timescales ranging from a few tens of seconds down to a few tens of milliseconds can arise from all radii, and subsequently propagate inward toward the black hole through the accretion flow (Lyubarskii1997; Kotov2001). At each radius, fluctuations are expected to propagate most efficiently at frequencies comparable to the local viscous timescale. Consequently, variability generated in the inner regions of the flow occurs at higher frequencies than that arising from larger radii. The fluctuation at each radius should be viewed as the superposition of the local fluctuation and those propagated from the outer region. This propagating fluctuation model naturally explains the observed correlations between variability in the soft and hard X-ray bands (Kotov2001; Zhan2025).
@@
In this work, the spectra of the coherent variability were well fitted with the model Tbabs*(diskbb+Nthcomp). It suggests that the BBN variability, on both long and short timescales, is not dominated by a single spectral component but rather by a combination of a disk and a Comptonization component. The modeling of the covariance spectra aligns with our interpretation in the previous section, indicating that the coherent variability at high energies is predominantly attributed to Comptonization of disk seed photons. Within the propagating-fluctuation model, the short-timescale variability arises from the inner region of the accretion flow, while long-timescale variability originates from larger radii (Kotov2001). In this framework, one would expect the Comptonization component of the variability spectra to exhibit progressively higher electron temperatures at higher Fourier frequencies (axelsson2018). However, we found that the Comptonization component on short timescales has a lower electron temperature than that on long timescales. This may indicate that the seed photon associated with Comptonization on the short timescale has a lower temperature than those linked to the long timescale.
@@
To understand the different temperatures of the disk seed photons inferred from the variability spectra, we consider two Comptonization continua. Specifically, the variability on short timescales originates in a central Comptonization region with a large scale height, while the long-timescale variability comes from a region located at somewhat larger radii. Zdziarski_2021a suggested that emission from the elevated Comptonization region is reflected off the disk at relatively large radii. Similarly, we propose that the Comptonization region with a large scale height primarily interacts with the outer disk at larger radii, whereas the lower Comptonization region interacts with the adjacent inner disk. Consequently, the Comptonized disk emission from the elevated region has a lower temperature, leading to the softer variability observed on short timescales.
@@
Within this framework, if the central Comptonization region has a larger vertical extent, it is expected to be illuminated by disk photons originating from larger radii, which have lower temperatures. This results in the Comptonization emission in this region having a lower electron temperature. Therefore, we could interpret the evolution of the electron temperature on different timescales, presented in Fig. 7, in terms of structural changes in the central Comptonization region. At the beginning of the hard state, the two Comptonization regions have nearly the same vertical height, making their covariance spectra indistinguishable. During the rise phase of the hard state (MJD 58192 to 58205), the central region extends vertically, increasing its height. As a result, the electron temperature associated with the short-timescale variability decreases, which corresponds to the deficit of hard X-ray variability on the short timescale (see the right panel of Fig. 6). During the decay of the hard state (after MJD 58205), as the central region contracts and its spatial extent decreases, the electron temperature increases on short timescales. Ultimately, after roughly MJD 58240, the two regions reach comparable heights, resulting in spectral properties that are indistinguishable once again.
5 Acknowledgements

### v2

@@ III.1.3 The root mean square and covariance spectra
The time-averaged X-ray spectra alone do not capture how the individual spectral components vary relative to each other on timescales of seconds to minutes. Instead, the PDS and coherence analyses reveal significant differences in variability across various energies and frequencies (Paice et al., 2021; Tetarenko et al., 2019, 2021). To further investigate the properties of the variability and their correlations, energy-dependent spectral-timing analyses are essential.
@@
Wilkinson and Uttley (2009) developed a technique known as the covariance spectrum. This covariance technique can be considered a matched filter applied to the data. By utilizing the PDS of both the reference band and the band of interest, we can derive the cross-spectrum and subsequently obtain a Fourier frequency-resolved covariance spectrum. This approach allows us to identify the variability in the energy channel being examined that is correlated with the reference band (Uttley et al., 2014). The covariance spectrum is closely related to the RMS spectrum. The covariance can be expressed as:
$Cv\left(\nu_{j}\right)=\langle x\rangle\sqrt{\gamma^{2}\left(\nu_{j}\right)}\sqrt{\left(\bar{P}_{X}\left(\nu_{j}\right)-P_{X,\text{noise}}\right)\Delta\nu_{j}},$ (1)
where $\gamma^{2}$ represents the coherence between two light curves within the specified frequency range. The term $\sigma_{\rm X}(\nu_{j})\equiv\sqrt{\left(\bar{P}_{X}\left(\nu_{j}\right)-P_{X,\text{noise}}\right)\Delta\nu_{j}}$ indicates the fractional RMS of the energy channel within that frequency range. It is clear that when the coherence is unity, the covariance and RMS are identical. For further details, please refer to the Appendix.
@@
We can obtain the RMS spectrum by integrating the PDS over a specified frequency range. The RMS is background-corrected using the formula $RMS=\sqrt{P}(S+B)/S$, where S and B represent the source and background count rates, respectively, and P is the power calculated over the selected frequency (Bu et al., 2015). The RMS spectrum illustrates the overall intrinsic variability, while the covariance spectrum emphasizes the variability that correlates with the reference band. If there is a single variability pattern across the entire energy band and coherence is perfect (unity), the shape of the covariance spectrum will match that of the RMS spectrum (Wilkinson and Uttley, 2009).
@@
Figure 4: Fractional covariance and RMS spectra. The top panel shows the fractional covariance (red circles) and fractional RMS (red squares) spectra on the short timescale. The bottom panel shows the corresponding spectra on the long timescale (blue circles and blue squares, respectively). The 2–10 keV band is selected as the reference.
III.1.4 Fractional covariance spectra
@@
By dividing the covariance spectra by the time-averaged spectra, we derive the fractional covariance spectra. Similar to the fractional RMS spectra, the fractional covariance spectra provide insight into the relative contributions of different components to the observed variability.(Cassatella et al., 2012).
@@
Figure 4 displays the fractional covariance and RMS spectra obtained on both long and short timescales, using the 2-10 keV reference band. At low energies, the fractional RMS spectra and the fractional covariance spectra closely align, appearing almost flat. And the coherence approaches unity. In contrast, at high energies, the fractional covariance spectra exhibit a power-law slope. Notably, the fractional covariance spectrum over the long timescale is significantly harder than that over the short timescale.
III.1.5 Covariance ratio
@@
We calculate covariance ratios by dividing the long-timescale covariance spectrum by the short-timescale covariance spectrum. By utilizing more than 70 observations from Insight-HXMT during the X-ray hard state, we investigated the evolution of the covariance ratio throughout the outburst. As mentioned earlier, due to the enormity of the dataset, here we only display the covariance ratios for the five representative epochs of observations in Figure 5. The covariance ratios remain relatively constant with energy below approximately 30 keV. However, above 30 keV, the covariance ratios increase with energy, indicating a hard excess of variability on long timescales. In other words, the trend of coherence toward higher energies on long timescales differs from that on short timescales. This suggests that the variability in the hard band exhibits distinct behaviors on both long and short timescales. By analyzing the significant changes in the covariance ratio with energy, we can see in a model-independent way that the variability of the hard band (>> 30 keV) emission, which is linearly correlated with the variability of the soft band (2-10 keV) emission, exhibits distinct behaviors on both long and short timescales. An increase in the covariance ratio may indicate additional Comptonization variability on long timescales.
A clear evolution in the covariance ratio is apparent across different epochs (Figure 5). In the first epoch (MJD 58194), the covariance ratio increases gradually with energy but displays a relatively shallow slope. In subsequent epochs, this energy-dependent trend becomes more pronounced. On MJD 58203, near the peak of the long-term light curve, the covariance ratio exhibits a significantly steeper slope than in other observations. After MJD 58203, the slope gradually decreases. This evolution indicates distinct X-ray variability behavior during the rising phase (before MJD 58200) and the decaying phase (after MJD 58200).
@@
III.2 Modeling of the covariance spectra
@@
In this study, we examined the covariance spectra, which are related to spectral variability. We have observed a significant hard excess of the variability over long timescales. To identify the origin of the hard excess, in this section, we try to fit the long- and short-timescale covariance spectra.
@@
For the covariance spectra, we employed a simple model Tbabs*(diskbb+Nthcomp), as the reflection features are not required due to the limited energy resolution of the covariance spectra. For each observation, we simultaneously fit the short- and long-timescale covariance spectra while constraining the absorbing column density of Tbabs to $N_{\mathrm{H}}=1.5\times 10^{21}\rm cm^{-2}$, which corresponds to the value used in the fits of the time-averaged energy spectra (You et al., 2021). The remaining parameters were allowed to vary freely between the covariance spectra. 72 of 76 covariance spectra can be fitted well with the reduced $\chi^{2}<1.0$.
The best-fitting covariance spectra and the time-averaged spectrum are presented in Figure 6. Part of the fitted parameters for energy and covariance spectra are listed in Table 2. The covariance spectra on different timescales are consistent at low energies but show different feature on at high energies, which implies that the hard excess is not associated with the soft component, but instead arises entirely from the hard X-ray emission.
Assuming that the variability of the Compton component mainly results from changes in its normalization, the covariance spectrum in its dominant energy band would closely mirror the shape of the time-averaged energy spectrum. The best-fitting parameters for the short-timescale covariance spectrum are generally consistent with those from the long-timescale covariance spectrum, except for the electron temperature. Interestingly, the electron temperature measured on the long timescale is significantly higher than that measured over the short timescale.
@@
We examine the temporal evolution of the spectral fit parameters. During the early hard state, the long-timescale electron temperature exceeds the short-timescale value (Figure 7). The long-timescale temperature remains relatively constant throughout the hard state, while the short-timescale temperature first decreases then rises, displaying an anticorrelation with X-ray luminosity. The photon index remains consistent between both variability components throughout the hard state. We estimate the optical depth of the Comptonizing component using Equation (2) from Yan et al. (2020). The optical depth exhibits markedly different temporal evolution on short versus long timescales (Figure 7).
Figure 7: Temporal evolution of the best-fitting parameters of the covariance spectra on long (blue dots) and short (red dots) timescales. From top to bottom, the panels show the evolution of the electron temperature (in keV), the photon index $\Gamma$, and the optical depth $\tau$ of the corona, respectively.
IV Discussion
In this study, we study energy-dependent coherence up to 150 keV and discover that the hard X-ray variability above approximately 30 keV becomes increasingly incoherent with the soft X-ray variability (2-10 keV) (see Fig. 3). This finding suggests that at least two distinct Comptonization components contribute to the overall X-ray variability. One component varies linearly with the soft X-ray variability and primarily emits below 30 keV. The other component varies incoherently with the soft X-ray variability and primarily emits above 30 keV. Hereafter, we refer to these components as the coherent component and the incoherent component, respectively.
@@ IV Discussion
Furthermore, we fit the covariance spectra for both timescales, using the model Tbabs*(diskbb+Nthcomp), to quantitatively characterize the covariance spectra (see Fig. 5). The parameters obtained from the short-timescale covariance spectrum are generally consistent with those from the long-timescale covariance spectrum, except for the electron temperature. Specifically, the electron temperature measured on the long timescale is significantly higher than that measured over the short timescale. Moreover, the evolution of electron temperature differs notably between the two timescales (see Fig. 7). The electron temperature associated with long-timescale variability is higher than that associated with short-timescale variability, indicating a deficit in hard X-ray variability on the short timescale. Specifically, the electron temperature for the long-timescale covariance spectrum remains relatively constant throughout the hard state, while the short-timescale electron temperature initially decreases before exhibiting a distinct increase. Thus, the evolution of the electron temperature for short timescale variability directly influences the evolution of the covariance ratio. In this case, we suggests that the increase in covariance ratios above approximately 30 keV, corresponding to the hard excess on long timescales, is caused by a deficit of hard X-ray variability on short timescales.
IV.1 On the incoherent variability: its origin
@@
The variability of Comptonization emission is influenced by variability of both the seed photons and electrons. Thus, the observed incoherent variability is likely a consequence of the complexity of these two components. In this section, we will discuss the potential regions where Comptonization occurs and the origins of the seed photons in order to better understand the nature of this variability. Furthermore, we will explore the association between the Comptonizing source and the flows near the BH.
IV.1.1 Multiple Comptonization emission
@@
One possibility accounting for the observed incoherent variability is that multiple Comptonization emissions, as proposed in previous studies, are involved. For instance, Dziełak et al. (2021) studied the BBN of the PDS for MAXI J1820+070. They found that the BBN could be modeled with four Lorentzian components, each exhibiting significant variability. At least two Comptonization components, characterized by different electron temperatures and optical depths, are required to accurately fit both the variability and time-averaged spectra. Zdziarski et al. (2021) studied the broadband spectra and found that a multi-Comptonization model can fit the spectra well. In their scenario, the system consists of two distinct Comptonization components: a harder component originating from a larger-scale height accretion flow located downstream of the truncation radius, and a softer component forming a corona over the inner disk. These separate regions produce distinct Compton reflection signatures. Wang and Zhang (2024) examined the evolution of the frequency-dependent RMS-flux relation in MAXI J1820+070 using Insight-HXMT data. The timing features and their evolution support the existence of two distinct Comptonization regions: one that remains radially extended, allowing fluctuations to propagate from the disk to the corona, and another that is vertically extended and continues to expand during the later stages of the initial hard state.
@@
It should be noted that we cannot completely rule out a single source of Comptonization. The exponential cut-off function is highly nonlinear(Uttley and Malzac, 2024), such that variations in the cut-off energy could also lead to a reduction in coherence. If the variability of the Comptonizing component arises not only from changes in its normalization but also from variations in the cut-off energy, the resulting variability spectrum may differ from the time-averaged spectrum. Owing to the non-linear nature of the cut-off function, such spectral variability can also lead to a decrease in coherence. It seems plausible that the incoherent variability could be partially attributed to this type of non-linear spectral variability. However, to draw more definitive conclusions, we would need further quantitative modeling that might suggest the presence of distinctly varying components beyond mere non-linearity. In this work, we assume that the variability of the Comptonizing component is primarily driven by changes in its normalization.
IV.1.2 Multiple fields of seed photon
@@
Another possibility responsible for the incoherent variability is associated with the complexity of the seed photon field. To examine this, we first consider disk photons as the seed photons involved in the inverse Compton scattering process. Uttley and Malzac (2024) constructs a model in which the interplay between the disk seed photons and the corona is taken into account to simulate the Comptonization variability. Based on their model, we conducted a simulation considering 10 groups of disk photons with different seed energies. All of these photons were Compton up-scattered by thermal electrons in the corona via inverse-Compton processes. Then we derive the overall variability spectrum over a broad energy band. Since the BBN is thought to arise from propagating accretion rate fluctuations, the input seed photon variations of all 10 groups are expected to be almost identical. In this case, we found that we could not reproduce the reduction of the coherence in the hard X-ray band (see Fig. 3), as all the resulting power-law emission variations remained coherent. However, when the input seed photon variations of the 10 groups are assumed to be independent of each other, a broad reduction in coherence naturally emerges. This suggests that the seed photons may originate from multiple components, rather than exclusively from the accretion disk.
@@
In this framework, we conclude that the seed photons for inverse-Compton scattering should have two origins: the disk blackbody and synchrotron emission (Yang et al., 2015), and it is the Comptonization of synchrotron emission as seed photons that is responsible for the incoherent variability.
IV.2 On the coherent variability: electron temperature on different timescales
@@
Variability on timescales ranging from a few tens of seconds down to a few tens of milliseconds can arise from all radii, and subsequently propagate inward toward the black hole through the accretion flow (Lyubarskii, 1997; Kotov et al., 2001). At each radius, fluctuations are expected to propagate most efficiently at frequencies comparable to the local viscous timescale. Consequently, variability generated in the inner regions of the flow occurs at higher frequencies than that arising from larger radii. The fluctuation at each radius should be viewed as the superposition of the local fluctuation and those propagated from the outer region. This propagating fluctuation model naturally explains the observed correlations between variability in the soft and hard X-ray bands (Kotov et al., 2001; Zhan et al., 2025).
@@
In this work, the spectra of the coherent variability were well fitted with the model Tbabs*(diskbb+Nthcomp). It suggests that the BBN variability, on both long and short timescales, is not dominated by a single spectral component but rather by a combination of a disk and a Comptonization component. The modeling of the covariance spectra aligns with our interpretation in the previous section, indicating that the coherent variability is predominantly attributed to Comptonization of disk seed photons. Within the propagating-fluctuation model, the short-timescale variability arises from the inner region of the accretion flow, while long-timescale variability originates from larger radii (Kotov et al., 2001). In this framework, one would expect the Comptonization component of the variability spectra to exhibit progressively higher electron temperatures at higher Fourier frequencies (Axelsson and Done, 2018). However, we found that the Comptonization component on short timescales has a lower electron temperature than that on long timescales. This may indicate that the seed photon associated with Comptonization on the short timescale has a lower temperature than those linked to the long timescale.
@@
To understand the different temperatures of the disk seed photons inferred from the variability spectra, we consider two Comptonization continua. Specifically, the variability on short timescales originates in a central Comptonization region with a large scale height, while the long-timescale variability comes from a region located at somewhat larger radii. Zdziarski et al. (2021) suggested that emission from the elevated Comptonization region is reflected off the disk at relatively large radii. Similarly, we propose that the Comptonization region with a large scale height primarily interacts with the outer disk at larger radii, whereas the lower Comptonization region interacts with the adjacent inner disk. Consequently, the Comptonized disk emission from the elevated region has a lower temperature, leading to the softer variability observed on short timescales.
@@
Within this framework, if the central Comptonization region has a larger vertical extent, it is expected to be illuminated by disk photons originating from larger radii, which have lower temperatures. This results in the Comptonization emission in this region having a lower electron temperature. Therefore, we could interpret the evolution of the electron temperature on different timescales, presented in Fig. 7, in terms of structural changes in the central Comptonization region. At the beginning of the hard state, the two Comptonization regions have nearly the same vertical height, making their covariance spectra indistinguishable. During the rise phase of the hard state (MJD 58192 to 58200), the central region extends vertically, increasing its height. As a result, the electron temperature associated with the short-timescale variability decreases, which corresponds to the deficit of hard X-ray variability on the short timescale (see the right panel of Fig. 6). During the decay of the hard state (after MJD 58200), as the central region contracts and its spatial extent decreases, the electron temperature increases on short timescales. Ultimately, after roughly MJD 58240, the two regions reach comparable heights, resulting in spectral properties that are indistinguishable once again.
V Acknowledgements